A fraction becomes 9/11 , if 2 is added to both the numerator and the denominator. If, 3 is added to both the numerator and the denominator it becomes 5/6. Find the fraction.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Here is the required solution:

Let the fraction be x/y.
According to the question,
(x+2) /(y+2) = 9/11
11x + 22 = 9y + 18
11x – 9y = -4 …………….. (1)
(x+3) /(y+3) = 5/6
6x + 18 = 5y +15
6x – 5y = -3 ………………. (2)
From (1),
we get
x = (-4+9y)/11 …………….. (3)
Substituting the value of x in (2),
we get
6(-4+9y)/11 -5y = -3
-24 + 54y – 55y = -33
-y = -9
y = 9 ………………… (4)
Substituting the value of y in (3),
we get
x = (-4+9×9 )/11 = 7
Hence the fraction is 7/9.

Thanks